OpenStudy (anonymous):

whale traveled 4 miles per hour, covering approximately 100 miles each day of his migration. What is the Whales speed of travel in kilometers per hour?

OpenStudy (john_es):

\[4 \frac{\text{miles}}{\text{hour}}\cdot\frac{1.609 \ \text{kilometers}}{1\ \ \text{mile}}=6.436\ km/h\]

OpenStudy (john_es):

The point is decimal point.

OpenStudy (anonymous):

thank you so much